On January 1, 2018, Engineers Credit Union (ECU) issued 8%, 20-year bonds payable with face value of $900,000. The bonds pay interest on June 30 and December 31 Read the requirements Requirement 1 . If the market interest rate is 7% when ECU issues its bonds, will the bonds be priced at face value, at a premium, or at a discount? Explain The 8% bonds issued when the market interest rate is 7% will be priced at la premium . They are attractive | in this market, so investors will pay more than face value to acquire them. Requirement 2. If the market interest rate is 9% when ECU issues its bonds, will the bonds be priced at face value, at a premium, or at a discount? Explain The 8% bonds issued when the market interest rate is 9% will be priced at la discount I. They are | unattractive | in this market, so investors will pay less than face value to acquire them. Requirement 3. The issue price of the bonds is 98. Journalize the bond transactions. (Assume bonds payable are amortized using the straight-line amortization method. Record debits first, then credits. Select explanations on the last line of the journal entry. Round your answers to the nearest whole dollar.) a. Journalize the issuance of the bonds on January 1, 2018 Debit Date 2018 Jan. 1 Accounts and Explanation Credit b. Journalize the payment of interest and amortization on June 30, 2018 Accounts and Explanation Date 2018 Jun. 30 Debit Credit c. Journalize the payment of interest and amortization on December 31, 2018 Accounts and Explanation Date 2018 Dec. 31 Debit Credit d. Retirement of the bond at maturity on December 31, 2037, assuming the last interest payment has already been recorded Date Accounts and Explanation Debit Credit 2037 Dec.
Answer #1
3
Debit Credit
2018
Jan. 1 Cash 882000 =900000*0.98
Discount on Bonds payable 18000
       Bonds payable 900000
Issued bonds at a discount
Jun. 30 Interest expense 36450
       Discount on Bonds payable 450 =18000/40
       Cash 36000 =900000*8%/2
Paid semiannual interest and amortized discount
Dec 31 Interest expense 36450
       Discount on Bonds payable 450 =18000/40
       Cash 36000 =900000*8%/2
Paid semiannual interest and amortized discount
2037
Dec 31 Bonds payable 900000
       Cash 900000
Retired Bonds at maturity